# DailyDVS-200: A Comprehensive Benchmark Dataset for Event-Based Action Recognition

### v1

@@ 1 Introduction
Although some benchmark datasets have been proposed, many of them are synthetic datasets captured by pointing event cameras at screens displaying RGB images [40], or by converting commonly used RGB action recognition datasets into simulated event streams [4, 23, 50, 21]. However, simulated datasets often result in information loss due to their multi-stage nature, making it difficult to achieve the effectiveness of real event datasets in practice. While several large-scale real-world benchmark event datasets have been proposed recently [1, 4, 49, 36, 33, 56, 17, 13], such as DvsGesture [1], PAF [36], and Hardvs [56], they are limited by small scale, few categories, and limited diversity of individuals. Additionally, $\mathrm{THU}^{\mathrm{E\text{-}ACT}}$-50 [17] consists only of simple actions in fixed scenes, and although $\mathrm{THU}^{\mathrm{E\text{-}ACT}}$-50-CHL [17] introduces challenging actions, it is still limited by the number and diversity of actions. Bullying10K [13] focuses exclusively on 10 types of bullying actions, making it difficult to demonstrate the superiority of models. Therefore, there is an urgent need for a comprehensive dataset that takes into account various factors to address these issues.
@@ 4 Experimental Results
In this section, we employ various methods to conduct cross-subject and multi-group evaluations on our dataset. Firstly, we test our dataset using multiple approaches, including frame-based, learnable-based, token-based, and spike-based methods. Secondly, within the frame-based approach, we explore the impact of frame gap and time step for generating event frames on the experiment results of our dataset. Additionally, we demonstrate the performance of multiple groups in various event-based action recognition models. Finally, we validate the challenge of our dataset by applying the same experimental settings to an existing large-scale action recognition dataset.
@@ 4.1.1 Evaluation of different methods
4.1.2 Evaluation of different groups
Figure 5: Evaluation of using different sizes of Moving camera set for action recognition.
To mitigate the bias in testing groups to camera motion attributes, we conducted training and testing on the TSM[29] model using different proportions of dynamic and static data. Our experimental results, as illustrated in Fig. 5, show a noticeable decrease in the accuracy of action evaluated under both dynamic and static cameras as the proportion of dynamic data increases. Contrary to expectations, there was no improvement in accuracy with the increasing volume of event data captured by dynamic cameras. This suggests that current action recognition models struggle to effectively learn action data captured by dynamic cameras, highlighting an urgent issue that requires resolution.
4.1.3 Evaluation of different frame settings
@@ 4.1.4 Detailed analysis to actions and methods
We also conducted a detailed analysis of the performance of different action recognition methods on our proposed DailyDVS-200 dataset. We took four best-performing methods (SlowFast[16], Swin-T[34], EST[18], Spikformer[59]) as examples for analysis. Firstly, we plot the confusion matrices of these methods. The confusion matrix of Swin-T[34] is shown in Fig. 6 as an example, we can see there are still many actions cannot be correctly classified. Specifically, considering the large number of action categories, we also analyzed the top-10 accurate actions and top-10 incorrect actions for each method (see Table 5).
@@
4.2 Comparison with other large-scale datasets
To demonstrate the diversity and challenging nature of our proposed dataset, we retrained and tested some of the popular models on the current large-scale event datasets [17, 56, 13] under the same training configuration. In this section, we used the framed-based model TSM[29] and the recently proposed baseline method ESTF in Hardvs[56]. Firstly, we converted all the datasets into frame sequences with a time interval of 0.5s and conducted training from scratch. Due to the absence of subject partitions in some datasets, we randomly split the datasets in a 6:1:3 ratio for experimentation. Our experimental results are shown in Table 6. It can be noted that these models has achieved the lowest accuracy on our dataset, with a top-1 accuracy of 31.29% only, indicating that our proposed DailyDVS-200 is currently the most challenging large-scale dataset. Meanwhile, we further evaluated TSM [29] with Kinetics-400[19] pretraining and found that these models still performed effectively under the event-based frame sequence, with DailyDVS-200 remaining the most challenging dataset, achieving a top-1 accuracy of 65.9%.
Table 6: Comparison of our proposed dataset with other large-scale datasets. represent the results pretrained on Kinetics-400[19].
Methods TSM[29] ESTF[56]
$\mathrm{THU}^{\mathrm{E\text{-}ACT}}$-50[17] 95.60 / $98.75^{{\dagger}}$ 95.25
$\mathrm{THU}^{\mathrm{E\text{-}ACT}}$-50-CHL[17] 49.07 / $83.83^{{\dagger}}$ 49.50
Hardvs[56] 97.33 / $98.55^{{\dagger}}$ 96.67
Bullying10K[13] 74.22 / $91.90^{{\dagger}}$ 84.72
DailyDVS-200(Ours) 36.05 / $\textbf{65.90}^{{\dagger}}$ 31.29

### v2

@@ 1 Introduction
Although some benchmark datasets have been proposed, many of them are synthetic datasets captured by pointing event cameras at screens displaying RGB images [42], or by converting commonly used RGB action recognition datasets into simulated event streams [4, 24, 52, 22]. However, simulated datasets often result in information loss due to their multi-stage nature, making it difficult to achieve the effectiveness of real event datasets in practice. While several large-scale real-world benchmark event datasets have been proposed recently [1, 4, 51, 38, 35, 58, 29, 18, 13, 15], such as DvsGesture [1], PAF [38], and Hardvs [58], they are limited by small scale, few categories, and limited diversity of individuals. Additionally, $\mathrm{THU}^{\mathrm{E\text{-}ACT}}$-50 [18] consists only of simple actions in fixed scenes, and although $\mathrm{THU}^{\mathrm{E\text{-}ACT}}$-50-CHL [18] introduces challenging actions, it is still limited by the number and diversity of actions. Bullying10K [13] focuses exclusively on 10 types of bullying actions, making it difficult to demonstrate the superiority of models. Therefore, there is an urgent need for a comprehensive dataset that takes into account various factors to address these issues.
@@ 4 Experimental Results
In this section, we employ various methods to conduct cross-subject and multi-group evaluations on our dataset. Firstly, we test our dataset using multiple approaches, including frame-based, learnable-based, token-based, and spike-based methods. Secondly, within the frame-based approach, we explore the impact of frame gap and time step for generating event frames on the experiment results of our dataset. Additionally, we demonstrate the performance of multiple groups in various event-based action recognition models. Finally, we validate our dataset’s complexity using standard experimental setups from a large-scale action recognition dataset, with details and findings in the supplementary material.
@@ 4.1.2 Evaluation of different groups
To mitigate the bias in testing groups to camera motion attributes, we conducted training and testing on the TSM[31] model using different proportions of dynamic and static data. Our experimental results, as illustrated in Fig. 5 (a), show a noticeable decrease in the accuracy of action evaluated under both dynamic and static cameras as the proportion of dynamic data increases. Contrary to expectations, there was no improvement in accuracy with the increasing volume of event data captured by dynamic cameras. This suggests that current action recognition models struggle to effectively learn action data captured by dynamic cameras, highlighting an urgent issue that requires resolution.
4.1.3 Evaluation of different frame settings
@@
Figure 5: (a) Evaluation of using different sizes of Moving camera set for action recognition. (b) Confusion matrix of Swin-T[36].
@@ 4.1.4 Detailed analysis to actions and methods
We also conducted a detailed analysis of the performance of different action recognition methods on our proposed DailyDVS-200 dataset. We took four best-performing methods (SlowFast[17], Swin-T[36], EST[19], Spikformer[61]) as examples for analysis. Firstly, we plot the confusion matrices of these methods. The confusion matrix of Swin-T[36] is shown in Fig. 5 (b) as an example, we can see there are still many actions cannot be correctly classified. Specifically, considering the large number of action categories, we also analyzed the top-10 accurate actions and top-10 incorrect actions for each method (see Table 5).
@@ 5 Conclusion
Acknowledgements
This work was partly supported by the Chinese Defense Advance Research Program (50912020105).